OpenStudy (mony01):

Let vectors A⃗ =(1,0,−3), B⃗ =(−2,5,1), and C⃗ =(3,1,1). Calculate the following, expressing your answers as ordered triples (three comma-separated numbers). B⃗ ×C⃗

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Cross_product

OpenStudy (owlfred):

Do you know the general expression for a cross product in terms of components? It's a simple 3x3 determinant.

OpenStudy (mony01):

oh do i need to find the determinant?

OpenStudy (owlfred):

Yes, you need to evaluate the determinant. If you don't know what they are, never really mind - it's just another overwhelming word that is truly simple. In general, if \(\vec A = (x_1, y_1, z_1)\) and \(\vec B = (x_2, y_2, z_2)\), then... \(\vec A \times \vec B = \hat i (y_1 z_2 - y_2 z_1) - \hat j (x_1 z_2 - x_2z_1) + \hat k (x_1 y_2 - x_2 y_1)\)

OpenStudy (owlfred):

In your notation, it is \((y_1 z_2 - y_2 z_1 , ~x_2 z_1 - x_1 z_2, ~x_1 y_2 - x_2 y _ 1) \)

OpenStudy (mony01):

what if it was to find A⃗ ×(B⃗ ×C⃗ )

OpenStudy (owlfred):

Then you would first find \(\vec B \times\vec C\), then cross that with \(\vec A\). Careful, though, with the order - cross product is not commutative.

OpenStudy (mony01):

so BxC=(4,5,-17) and A=(1,0,-3) so i do the same thing as above

Parth (parthkohli):

Yes.

OpenStudy (mony01):

thanks

Parth (parthkohli):

No problem!